OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4x and y = 2x + 2 intersect are the solutions of the equation 4x = 2x + 2. Part B: Make tables to find the solution to 4x = 2x + 2. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4x = 2x + 2 graphically?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

i already have part a

OpenStudy (the_fizicx99):

Well for B then just make a table, plug in -3 for x and go increasing till you get to +3

OpenStudy (anonymous):

that doesnt help how am i suppose to do the table like ....

OpenStudy (ranga):

x (4x) (2x+2) -3 -12 -4 -2 -8 -2 ....

OpenStudy (anonymous):

oh

OpenStudy (ranga):

x should go: -3, -2, -1, 0, 1, 2, 3

OpenStudy (anonymous):

and i keep going from there .. and do the same table but with a positive 3 as well right.. it would justbe the same thing though except without negatives...

OpenStudy (the_fizicx99):

4(-3) = 2(-3) + 2 -12 = -6 + 2 -12 = -4 _______________ 4(-2) = 2(-2) + 2 -8 = -4 + 2 -8 = -2 _________________ 4(-1) = 2(-1) + 2 -4 = -2 + 2 _________________ 4(0) = 2(0) + 2 0 = 2 __________________ 4(1) = 2(1) + 2 4 = 4 ____________________ 4(2) = 2(2) + 2 8 = 4 + 2 8 = 6 _____________________ 4(3) = 2(3) + 2 12 = 6 + 2 12 = 8 _____________________

OpenStudy (the_fizicx99):

At the x value of 1 they would intersect btw

OpenStudy (anonymous):

both of you gave me different table eamples

OpenStudy (the_fizicx99):

._. Really? lol, no we didn't

OpenStudy (the_fizicx99):

I just put " = "

OpenStudy (anonymous):

yeah cuz @ranga told me it should go like -3, -2, -1, 0, 1, 2, 3

OpenStudy (the_fizicx99):

.____________. Thats what I did, I plugged in those values for x

OpenStudy (anonymous):

dont be rude, not eveyone can be the Socrates of algebra

OpenStudy (ranga):

Both are the same. I showed the table entries in compact form and tHe_FiZiCx99 showed you the details of the calculation.

OpenStudy (the_fizicx99):

How was a rude?!? ;~;

OpenStudy (the_fizicx99):

I*

OpenStudy (ranga):

x (4x) (2x+2) -3 -12 -4 -2 -8 -2 -1 -4 0 0 0 2 1 4 4 2 8 6 3 12 8

OpenStudy (anonymous):

ok this is the table i made x (4x) (2x+2) -3 -12 -4 -2 -8 -2 -1 -4 0 0 0 2 1 4 4 2 8 6 3 12 8

OpenStudy (ranga):

when x = 1, both 4x and (2x+2) yield the same value. So x = 1 is the solution.

OpenStudy (anonymous):

omg we got the same one, i feel smarter now, u couldve saved me 2 minutes though haha

OpenStudy (the_fizicx99):

^^

OpenStudy (anonymous):

ok and for part c should i just use desmos ?

OpenStudy (ranga):

brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@tHe_FiZiCx99 for part c should i just use desmos ??

OpenStudy (the_fizicx99):

To graph it you would need to have a " y = "

OpenStudy (anonymous):

ughhhhhh and how do i do that -____________________- >.<

OpenStudy (the_fizicx99):

I'm kinda eh on that, though it says " y= 4x" but I'm not entirely sure :l hmm

OpenStudy (the_fizicx99):

Got it :DD

OpenStudy (ranga):

Use desmos or a graphing calculator if you have one, or just a simple graph paper to graph these two lines. Graph: y = 4x and y =2x + 2 Both will be straight lines. The point where they intersect will be the solution.

OpenStudy (the_fizicx99):

y = 4x y = 2x + 2 Plot that into desmos and you'll get (1,4) which is the solution 4 = 4(1) 4 = 2(1) + 2 4 = 4

OpenStudy (the_fizicx99):

They intersect at (1,4)

OpenStudy (ranga):

Yes they should intersect at (1,4) as we already determined from the table where when x = 1, both functions gave us 4.

OpenStudy (anonymous):

thats it for part C ?

OpenStudy (the_fizicx99):

Yes.

OpenStudy (anonymous):

thank you guys so much, can u help me on my new question

OpenStudy (the_fizicx99):

I have to get off, cya